OpenStudy (anonymous):

one number is eight more than twice another. If their sum is decreased by ten, the result is thirteen. Find the numbers.

OpenStudy (anonymous):

\[a=2b+8\]\[a+b-10=13=>a+b=23=>a=23-b\]\[=>a=23-b=2b+8=>3b+8=23=>3b=15=>b=5\]\[=>a+b=a+5=23=>a=18\]